OpenStudy (anonymous):

Suppose a,b, and c are positive integers such that a+b+c+ab+bc+ca+abc=1000. Find a+b+c.

OpenStudy (asnaseer):

You have posted the same question again. same hint applies - try to factor the expression on the left

OpenStudy (anonymous):

i dont know how to do that.

OpenStudy (asnaseer):

and make an attempt to factor it - you will soon see a pattern evolve

OpenStudy (asnaseer):

*at least make an attempt

OpenStudy (asnaseer):

I can start you off...

OpenStudy (asnaseer):

a + b + c + ab + bc + ca + abc = 1000 a(1+b) + b + c + bc + ca + abc = 1000 a(1+b) + c(1+b) + b + ca + abc = 1000 a(1+b) + c(1+b) + ca(1+b) + b = 1000 a(1+b) + c(1+b) + ca(1+b) + b + 1 - 1 = 1000 a(1+b) + c(1+b) + ca(1+b) + (1+b) = 1001 can you continue from here?

OpenStudy (anonymous):

let me try

OpenStudy (anonymous):

no idea

OpenStudy (asnaseer):

can you there is a common factor of (1+b) in all the terms on the left of the equals sign?

OpenStudy (anonymous):

yes

OpenStudy (asnaseer):

so first pull that common factor out - what do you end up with?

OpenStudy (asnaseer):

a(1+b) + c(1+b) + ca(1+b) + (1+b) = (1+b)( ... ) = 1001

OpenStudy (anonymous):

(1+b)( a+c+ac+1) = 1001

OpenStudy (anonymous):

then?

OpenStudy (asnaseer):

good - now notice the two terms that have a common factor of "c" - factor those next

OpenStudy (anonymous):

ok

OpenStudy (asnaseer):

what do you end up with?

OpenStudy (anonymous):

(1+a)(1+b)(1+c)=1001

OpenStudy (asnaseer):

perfect!

OpenStudy (asnaseer):

now notice the right hand side - split it into a product of prime numbers

OpenStudy (anonymous):

how?

OpenStudy (asnaseer):

find the factors of the number 1001

OpenStudy (asnaseer):

hint - try dividing it by 11

OpenStudy (anonymous):

91*11

OpenStudy (asnaseer):

now try factoring 91

OpenStudy (anonymous):

7*13

OpenStudy (asnaseer):

perfect!

OpenStudy (asnaseer):

so you ended up with: (1+a)(1+b)(1+c)=1001=7*11*13

OpenStudy (asnaseer):

since the product on the right involves primes numbers only, you can assert that each term on the left is equal to one of these primes

OpenStudy (anonymous):

oh i see thank you sir thank you very much

OpenStudy (asnaseer):

yw :)